how much of a 10% syrup solution should be added to 2 litres of a 30% syrup solution to get a 20% syrup solution
Found 2 solutions by stanbon, TimothyLamb:
Answer by stanbon(75887)

how much of a 10% syrup solution should be added to 2 litres of a 30% syrup solution to get a 20% syrup solution
--------------
Equation:
active + active = active
0.10x + 0.30*2 = 0.20(x+2)
-------
10x + 30*2 = 20x + 20*2
-----------------
10x = 10*2
----
x = 2 litres (amt. of 10% solution needed)
----------------------------------------------
Cheers,
Stan H.
============

Answer by TimothyLamb(4379)

10x + 30(2) = 20(x + 2)
10x + 60 = 20x + 40
10x = 20
---
answer
x = 2 liters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php